OpenStudy (anonymous):

What is the third–degree polynomial function such that f(0) = –24 and whose zeros are 1, 2, and 3?

OpenStudy (amistre64):

its really just a matter of setting up the product using the zeros as bait; then scaling it such that f(0) = -24

OpenStudy (anonymous):

to tell you the truth i really dont know

OpenStudy (amistre64):

given a root, the set up is such that: x - n = 0; when x=r; n=-r set up 3 products using the roots (x-1)(x-2)(x-3) now, determine your scalar by making x=0 a(-1)(-2)(-3) = -24 a(-6) = -24; when a=4 4(x-1)(x-2)(x-3) should do it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

well these are the possisble answers f(x) = 4x3 – 24x2 + 44x – 24 f(x) = 2x3 – 24x2 + 22x – 24 f(x) = 4x3 + 24x2 – 44x – 24 f(x) = 2x3 + 24x2 – 22x – 24

OpenStudy (amistre64):

can you narrow it down any?

OpenStudy (amistre64):

if anything, copy the solution i gave you into the wolf and see how it expands it ....